OpenStudy (anonymous):

An electron in a cathode-ray-tube (CRT) accelerates uniformly from 5.00 ✕ 104 m/s to 6.00 ✕ 106 m/s over 3.00 cm. (a) In what time interval does the electron travel this 3.00 cm? (b) What is its acceleration?

OpenStudy (irishboy123):

i would do these in reverse using the equation of motion \(v^2 = u^2 + 2ax\) will give you the acceleration, a, with least fuss using the equation of motion \(v = u + at\) will give you the time, t the 3 basic equations l are 1. \[v = u + at\] 2.\[x = ut + \frac{1}{2}at^2\] 3.\[v^2 = u^2 + 2ax\]